OpenStudy (anonymous):

6 – x – x = 18 (1 point) –6 6 9 –9

OpenStudy (geometry_hater):

you need to combine like terms then get x by itself

OpenStudy (anonymous):

6 - 2x = 18 -2x =18-6 -2x =12 x =12/-2 an u solve it?